The length of a rectangular hall is 5 metres more than its breadth. The area of the hall is 750 metre square. The length of the hall is-
Answer by laoman(51)

Given that the area of the hall is 750 m^2
We also know that the length is 5 m more than the breadth
If the length of the hall is l and the breadth is b
We can say that l = b+ 5 from the info given
Area = length * breadth
= (b+5) * b = 750
b^2 + 5*b = 750
b^2 + 5*b - 750 = 0
Solving the quadratic equation above,
Using factorization method, our factors are 30 and -25
A rule state that the negative inverse of the factors are the roots of a quadratic equation with 1 as the coefficient of the highest power
Therefore b = -30 or +25
But we cant have a negative breadth, therefore breadth, b = 25
Also, length,l = b+5 = 25+5=30
Therefore the length of the hall is 30 metres
Never forget the units
QED